A valószínűségszámítás dolgozat eredménye

1.  2.  3.  4.  5.  6.

diákok:

Ádám Zsolt  +  +  −  +  −  ±
Aradi Bernadett  +  +  +  +  +  ±
Balázs Éva  +  ±  −  +  ±  ±
Barna Timea  +  ±  −  −  +  +
Bócsi Judit  −  −  −  −  −  +
Bodnár Nelli  −  ±  −  +  +  +
Dobány Máté  +  ±  −  +  +  +
Fista Norbert  ±  +  +  +  +  +
Jenei Tamás  +  +  +  ±  −  +
Juhász Mónika  +  −  +  −  +  −
Karakas Éva  −  −  −  −  ±  −
Kész Timea  ±  −  −  −  ±  +
?Kiss Annamária  +  −  −  ±  ±  ±
Kovács György  −  −  −  −  −  +
Kovács Viktor  −  −  −  −  −  −
Kövér Ildikó  −  −  −  −  −  +
Léstár Adrienn  −  −  −  −  −  −
Lippai Zsolt  +  −  −  −  +  +
Mályi Nikoletta  −  −  −  −  −  +
Mezei Anita  −  ±  −  +  +  +
Mizsei Zsolt  −  ±  −  −  +  +
Nagy Beáta  +  −  −  +  +  ±
?Németh Ildikó  −  −  −  ±  −  −
Péter Antal  −  −  −  −  −  −
Pósa Csenge  −  ±  −  −  ±  +
Sikolya Kinga  +  +  +  +  +  +
Szilágyi Renáta  −  −  −  −  −  −
Szokol Patricia  +  +  +  +  +  +
Szoó Tibor  −  ±  +  +  ±  +
Tóth Andrea  −  −  −  −  −  ±
Tóth Eszter  −  −  −  −  +  ±
Tóth Veronika  +  −  −  −  ±  +
Tóth Zoltán  +  +  ±  −  +  +
Weinémer Andrea  ±  +  ±  −  −  +